OpenStudy (anonymous):

Is there an aid for remembering SOHCAHTOA and is there a mnemonic device for the SI system?

OpenStudy (anonymous):

SohCahToa IS the aid for remembering the basic trig functions. Where o is the side opposite the angle in question (theta), a is the side adjacent and h is the hypotenuse: \[\Large \begin{array}{l} \sin \theta = \frac{o}{h}\\ \cos \theta = \frac{a}{h}\\ \tan \theta = \frac{o}{a} \end{array}\]

OpenStudy (unklerhaukus):

What exactly do you mean the SI system, what are you trying to remember about it

OpenStudy (anonymous):

I can't think of any SI system mnemonics - just remember the base units: kilogram, metre, second, Ampere, Kelvin, candela and mole. I suppose you could arrange this in some kind of easy-to-recall mnemonic: k m s A K c m But I've always just remembered the units themselves.